OpenStudy (ashleyisakitty):

What is the difference in simplest form? n^2+3n+2/n^2+5n+6 - 2n/n+3

OpenStudy (austinl):

\(\displaystyle \bf \frac{n^2+3n+2}{n^2+5n+6}-\frac{2n}{n+3}\) This look correct Ashley?

OpenStudy (ashleyisakitty):

Yes thats correct @austinL

OpenStudy (mathstudent55):

Since you are subtracting fractions you need a common denominator. Start by factoring the numerator and denominator of the left fraction.

OpenStudy (bibby):

Here's a start, the numerator of the left piece factors into (n+2)(n+1)

OpenStudy (ashleyisakitty):

Right. and the denominator is (n+2)(n+3)

OpenStudy (bibby):

The n+2's cancel and you just subtract across

OpenStudy (ashleyisakitty):

so is it 1/n+3?

OpenStudy (bibby):

\[\frac{ (n+2)(n+1) }{ (n+3)(n+2) } = \frac{ (n+1) }{ (n+3) } \] \[\large \frac{ n+1 }{ n+3 }-\large \frac{ 2n }{ n+3 } = \frac{ n+1-2n }{ n+3 } = \frac{ n+1-2n }{ n+3 } =\frac{ 1-n }{ n+3 } \]

OpenStudy (bibby):

oops, there's a double in there

OpenStudy (ashleyisakitty):

Okay. I get it. Thanks @bibby and @mathstudent55.